# thanks: (No relevant financial or non-financial competing interests. Derived from preprint [23].)

### v3

@@ 14       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality [29]. Suppose that
@@
Note that we also need the conditions $\log\delta^{-1}>1/\varepsilon$ and
$s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log\delta^{-1}+1}{\varepsilon\log\delta^{-1}-1}$
@@
We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\varepsilon$ is equivalent to $t>\exp\left(\frac{1}{2\varepsilon}\right)$. Write
$l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\varepsilon}\right)\right\rfloor,\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\right\}.$ (30)
Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\varepsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
@@
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right)$
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log t^{2}\frac{8\log t^{2}+1}{\varepsilon\log t^{2}-1}\right).$
@@
$l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (31)
@@
For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (33)
A Monte Carlo simulation example based the Pareto distribution of type III (see Sec. 1.1) is shown in Fig. 1. Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e., $\chi$ defined in (9) becomes sufficiently small such that, in the $\sup$ expression of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is nonzero and the growth rate of the right-hand side with $\theta$ becomes small).
Figure 1: The performance of extended robust UCB
2.2 Further Analysis and Comparison of the Extended Robust UCB
Now we consider scenarios where specific values of $p$ and $q$ are chosen to further improve the efficiency of the extended robust UCB policy.
Case 1. Suppose that the $p^{\prime}$-th order of moments of reward distributions exist with $p^{\prime}>2$ for all arms. Then we can choose $p$, $q$ such that $p=2q\leq p^{\prime}$. The inequality (16) can be refined to
$\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]&=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2\mathbb{E}v_{q}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\\ &\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (34)
Hence, we can refine $C^{\prime}$ from (13) to
$C^{\prime}=\sqrt{12(C_{p,q}-1)},$ (35)
which can be much smaller than the original definition (13).
Let’s come back to the definition of $\chi_{M}$, which guarantees the three inequalities (26), (27) and (28) to hold. The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e., the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy may converge to the best arm faster.
Case 2. In addition to the assumption in Case 1, we add an assumption that the $4$-th order moments of reward distributions exist as in Lattimore [21]. Thus we can choose $p=4$ and $q=2$. In this case, our UCB in Definition 2 can be rewritten as
@@ 2.2 Further Analysis and Comparison of the Extended Robust UCB
In this case $C_{p,q}$ degenerates to the upper bound on kurtosis. The function $B(x)$ defined in (23) becomes
and the three inequalities (26), (27) and (28) are refined to
11 11 $>C^{\prime}\sqrt{x},$ (36)
11 11 $>48x\tau(x),$ (37)
$\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (38)
Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (33) will be significantly decreased. Now we compare the extended robust UCB and the scale free algorithm proposed in Lattimore [21], under the same assumptions specified at the beginning of Case 2.
We first address a flaw within Lemma 2 of Bubeck et al. [5]. This lemma, crucial in Lattimore’s analysis [21] for deriving the regret upper bound of the scale free algorithm, is examined more closely below. For clarity and coherence with the notations employed in this paper, we present the following adapted version.
A Flawed Statement [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
The last inequality in the proof of this statement in Bubeck et al. [5] assumes that $\exp(-k^{\prime}/8)\leq\delta$, and this is equivalent to $k^{\prime}\geq 8\log\delta^{-1}$. However, the definition of $k^{\prime}$ in this statement is
$k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor.$
To ensure that the inequality $\exp(-k^{\prime}/8)\leq\delta$ in their proof to hold, it is required that
$\lfloor s/2\rfloor\geq 8\log\delta^{-1},$ (39)
which is not assumed in the statement. As the statement is applied to the robust UCB, under the choice $\delta=t^{-2}$ and the definition $s:=s_{i}(t-1)$ as the number of times that arm $i$ was chosen before time $t$, Assumption (39) becomes
$\lfloor s_{i}(t-1)/2\rfloor\geq 16\log t.$
Hence the regret upper bound in Theorem 3 of Bubeck et al. [5] is at least $32\Delta_{i}\log T$.
Below we give the corrected version of the above statement. Again, we will adopt notations introduced in this paper for consistency.
Lemma 4.
Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$, where $s$ is chosen such that $\lfloor s/2\rfloor\geq 8\log\delta^{-1}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
Now we look into the results in Lattimore [21], whose Lemma 4 is a special version of the flawed Lemma 2 in Bubeck et al. [5]. Lattimore’s Lemmas 5 and 6 in [21] both use the flawed version to estimate the regret upper bound as in the proof of their Theorem 2. We do not intend to give details about fixing all of them, but it is important to point out that, after applying the correct version (i.e., adding Assumption (39)) as in Lemma 4, the upper bound estimation of $\mathbb{E}[s_{i}(T)]$ in Lattimore [21] is at least (the constants introduced in their paper are explicitly calculated with notations translated here):
$\begin{split}3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}\log({\rm e}^{1/8}\delta^{-1})+16\log\delta^{-1}.\end{split}$ (40)
The next theorem shows that for properly chosen parameter $\varepsilon>0$, the upper bound given in (33) is also tighter than (40) over long-run if $\Delta_{i}$ is small, i.e., arms are hard to be distinguished (the low-discrimination case).
Theorem 2.
Choose any $8<\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$ ($\approx 44.158$) and let $\delta=t^{-2}$. The upper bound given in (40) of $\mathbb{E}[s_{i}(T)]$ is greater than that given in (33) for sufficiently large $T$ and sufficiently small $\Delta_{i}$. That is, by dividing both the upper bound estimations by $\log\delta^{-1}$ and letting $T\to+\infty$, we have
$(8+\varepsilon)\left(\frac{8}{\varepsilon}+\max\left\{\frac{8}{\varepsilon},\frac{1}{\chi_{M}}\right\}\right)<3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}+16$ (41)
for sufficiently small $\Delta_{i}$. The regret upper bound of the scale free algorithm proposed in Lattimore [21] is thus larger than that of the extended robust UCB policy for $\varepsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
Proof.
@@
First note that, since $\Delta_{i}$ is sufficiently small and $\varepsilon$, $C^{\prime}$ are not related to $\Delta_{i}$, we only need to show (noticing that $\varepsilon>8$ implies $(8+\varepsilon)\frac{8}{\varepsilon}\leq 16$)
$(8+\varepsilon)\frac{1}{\chi_{M}}<3648\frac{v_{2}}{\Delta_{i}^{2}}.$ (42)
@@
$x^{\prime}:=\frac{(8+\varepsilon)\Delta_{i}^{2}}{3648v_{2}}<\chi_{M}$ (43)
for sufficiently small $\Delta_{i}$. In order to prove the above, we need to show that (36), (37) and (38) are all true for $x=x^{\prime}$ based on the definition of $\chi_{M}$. For (36), let $x=x^{\prime}$, we have $1>\frac{C^{\prime}\Delta_{i}}{8}\sqrt{\frac{\varepsilon+8}{57v_{2}}}$, which is true for small $\Delta_{i}$. Furthermore, (37) becomes
$1>6(8+\varepsilon)\Delta_{i}^{2}\frac{1}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(8+\varepsilon)v_{2}}},$
@@
$(\varepsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\varepsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\varepsilon+8)v_{2}}-6\Delta_{i}^{2}(\varepsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (44)
Taking the limit $\Delta_{i}\to 0$, (44) becomes $\left(6\sqrt{2}+9\right)(\varepsilon+8)<912$, which holds for $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. By continuity, we conclude that (43) holds for sufficiently small $\Delta_{i}$ if $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. ∎
Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound by only a constant factor and a constant term as in the scale free algorithm of Lattimore [21] (see (40)). First, we restate the lower bound derived in Lattimore [21] as follows.
Theorem 3.
Let $H_{\kappa_{0}}$ be the set of distributions that has kurtosis less than $\kappa_{0}$. Assuming $\Delta>0$ and $\kappa_{0}\geq 7/2$. Suppose that $X\in H_{\kappa_{0}}$ has a mean of $\mu$, a positive variance of $\sigma^{2}$, and a kurtosis of $k$. Then
[…]
Based on this, we can draw the following conclusion:
Theorem 4.
Let $\delta=t^{-2}$. Under case 2, the regret of the extended robust UCB policy applied on $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of the regret by a constant factor and a constant term.
Proof.

### v4

@@ Appendix E Proof of Lemma 3
by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality Rudin [29]. Suppose that
@@
Note that we also need the conditions $\log\delta^{-1}>1/\epsilon$ and
$s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log\delta^{-1}+1}{\epsilon\log\delta^{-1}-1}$
@@
We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
$l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\epsilon}\right)\right\rfloor,\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\right\}.$ (43)
Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
@@
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right)$
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\epsilon)\log t^{2}\frac{8\log t^{2}+1}{\epsilon\log t^{2}-1}\right).$
@@
$l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (44)
@@
For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
$\hfill\qed$
Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
Recall all three required assumptions which are related to $C_{p,q}$:
$1>$ $C^{\prime}x^{1-\frac{q}{p}},$ (47)
$1>$ $3\cdot 2^{2q}x^{q-1}\tau(x),$ (48)
$\Delta_{i}\geq$ $v_{q}^{1/q}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q}).$ (49)
From the previous deduction, we conclude that $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}$, where $l_{0}=\max\{\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}=O(\log T)$ and $l_{1}=\max\{l_{0},\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\rceil\}$. If we consider the case when replacing $C_{p,q}$ by $f(T)$ as long as T is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not rely on the index of a certain arm since those requirements are established for all arms to attain. Now we only need to figure out the order of $\frac{1}{\chi_{M}}$ to prove our statement.
For simplicity, we will discuss the three inequalities in turn. Recall that we use $f(T)$ instead of $C_{p,q}$ and focus on the case when $f(T)\geq C_{p,q}$, i.e. $T>t_{0}$ for some $t_{0}>0$.
@@ Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
$\frac{1}{\chi_{M}}=O(f(T)^{\frac{q}{p-q}}).$ (50)
[…]
$\chi_{M}\textless[\frac{1}{c_{4}\cdot\tau(\chi_{M})}]^{\frac{q}{q-1}}.$ (51)
After combining conditions given in (47) and (48), we see that (49) holds if
$c_{0}\geq c_{1}\cdot\chi_{M}^{1-\frac{1}{q}}\left[c_{2}+c_{3}(\frac{\tau(\chi_{M})}{1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})})^{\frac{1}{q}}\right].$
This indicates we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49). Simple analysis shows that left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$. Thus we only require that
$c_{0}\geq c_{0}c_{4}\chi_{M}^{q-1}\tau(\chi_{M})+c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$
$=\chi_{M}^{1-\frac{1}{q}}[c_{0}c_{4}\chi_{M}^{q+\frac{1}{q}-2}\tau(\chi_{M})+c_{1}c_{2}+c_{1}c_{3}\tau(\chi_{M})].$
By the fact that $\chi_{M}\leq 1$, the above inequality which is a stricter version of (49) holds if
$c_{0}\geq\chi_{M}^{1-\frac{1}{q}}[(c_{0}c_{4}+c_{1}c_{3})\tau(\chi_{M})+c_{1}c_{2}],$
i.e.
$\chi_{M}\leq[\frac{c_{0}}{k\cdot\tau(\chi_{M})+b}]^{\frac{q}{q-1}},$ (52)
where $k=c_{0}c_{4}+c_{1}c_{3},b=c_{1}c_{2}$.
Furthermore, by comparing (51) with (52), they simultaneously hold if
$\chi_{M}\leq\left[\frac{\min\{1,c_{0}\}}{\max\{k+b,c_{4}\}\cdot\tau(\chi_{M})}\right]^{\frac{q}{q-1}}=\left[d(1-(12(f(T)+1))^{\frac{q}{p}}\chi_{M}^{1-\frac{q}{p})}\right]^{\frac{q}{q-1}},$
where $d=\frac{\min\{1,c_{0}\}}{\max\{k+b,c_{4}\}}$. Since $\chi_{M}<1$, we only require that
$\chi_{M}\leq(\frac{d}{1+[12(f(T)+1)]^{\frac{q}{p}}d})^{\frac{1}{D}},~{}~{}~{}D=1-\max\{\frac{1}{q},\frac{q}{p}\}>0.$ (53)
Finally, through comparison between (50) and (53), we have $\frac{1}{\chi_{M}}=O(f(T)^{\max\{\frac{q}{pD},\frac{q}{p-q}\}})=O(f(T)^{\frac{q}{pD}})$. By setting $l_{0}=\max\{t_{0},\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}$, we have $l_{1}=O(f(T)^{\frac{q}{pD}}\log T)$ and $\mathcal{R}_{\pi}^{\mathcal{F}}(T)=O(l_{0}+l_{1}+\frac{5\pi^{2}}{6})=O(f(T)^{\frac{q}{pD}}\log T)$. $\hfill\qed$
Appendix G Proof of Theorem 2
@@
First note that, since $\Delta_{i}$ is sufficiently small and $\epsilon$, $C^{\prime}$ are not related to $\Delta_{i}$, we only need to show that (noticing that $\epsilon>8$ implies $(8+\epsilon)\frac{8}{\epsilon}\leq 16$)
$(8+\epsilon)\frac{1}{\chi_{M}}<3648\frac{v_{2}}{\Delta_{i}^{2}}.$ (54)
@@
$x^{\prime}:=\frac{(8+\epsilon)\Delta_{i}^{2}}{3648v_{2}}<\chi_{M}$ (55)
for sufficiently small $\Delta_{i}$. In order to prove the above, we need to show that (10), (11) and (12) are all true for $x=x^{\prime}$ based on the definition of $\chi_{M}$. For (10), let $x=x^{\prime}$, we have $1>\frac{C^{\prime}\Delta_{i}}{8}\sqrt{\frac{\epsilon+8}{57v_{2}}}$, which is true for small $\Delta_{i}$. Furthermore, (11) becomes
$1>6(8+\epsilon)\Delta_{i}^{2}\frac{1}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(8+\epsilon)v_{2}}},$
@@
$(\epsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\epsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\epsilon+8)v_{2}}-6\Delta_{i}^{2}(\epsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (56)
Taking the limit $\Delta_{i}\to 0$, (56) becomes $\left(6\sqrt{2}+9\right)(\epsilon+8)<912$, which holds for $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. By continuity, we conclude that (55) holds for sufficiently small $\Delta_{i}$ if $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. $\hfill\qed$
Appendix H Proof of Theorem 4